OpenStudy (brooke10239):

Order from least to greatest: 3 over 5 (3/5), 0.8, .06, 72%

jabez177 (jabez177):

0.8 is also knows as 8/10 because 8 is in the tenths place when you're reading a decimal.

jabez177 (jabez177):

So 8/10 is also knows as 4/5

jabez177 (jabez177):

And .06 did you mean 0.6 ?

OpenStudy (brooke10239):

Because you simplify, got it, and no it says .06

jabez177 (jabez177):

Okay. We will have the common denominator be 100 then.

OpenStudy (brooke10239):

Okay

jabez177 (jabez177):

So 3/5 is knows as 6/10 and 6/10 turns into 60/100 3/5 = 60/100 Can you do the next one? :)

jabez177 (jabez177):

The 8/10

OpenStudy (brooke10239):

80/100

jabez177 (jabez177):

Right. :) Now .06 6 is in the Hundreths place so can you guess what that will be in Fraction from? :)

OpenStudy (brooke10239):

6/100 ?

jabez177 (jabez177):

Yep! :) So far we have: 60/100 80/100 6/100 And now the 72%. We know that 72 percent means 72% of 100 and remember our keywords, of means a fraction so Can you try and figure out what 72% is as a fraction? :)

OpenStudy (brooke10239):

18/25 I think

jabez177 (jabez177):

Nope, it's 72/100 Any Percent sign you see, it is always out of 100

OpenStudy (brooke10239):

Oh, okay, I see

jabez177 (jabez177):

60/100 80/100 6/100 72/100 Can you list them from Least to Greatest now? :)

OpenStudy (brooke10239):

80/100, 72/100, 60/100, 6/100 or .8, 72%, 3/5, and .06

jabez177 (jabez177):

60/100 = 3/5 80/100 = 0.8 6/100 = 0.06 72/100 = 72% Now you have to list them from Smallest to Biggest.

jabez177 (jabez177):

Which 1 is the smallest first of all?

OpenStudy (brooke10239):

Oh yea, I did it backwards 0.06, 3/5, 72%, and 0.8

OpenStudy (brooke10239):

6/100 or 0.06

jabez177 (jabez177):

Bingo! You got it! Nice job! :)

OpenStudy (brooke10239):

Thanks for all the help!

jabez177 (jabez177):

You're welcome! :)